# Quantifying the robustness of the Auriga galaxy formation model

### v1

@@ Abstract
Numerical simulations have become an indispensable tool in astrophysics. To interpret their results, it is critical to understand their intrinsic variability, that is, how much the results change with numerical noise or inherent stochasticity of the physics model. We present a set of seven realisations of high-resolution cosmological zoom-in simulations of a Milky Way-like galaxy with the Auriga galaxy formation model. All realisations share the same initial conditions and code parameters, but draw different random numbers for the inherently stochastic parts of the model. We show that global galaxy properties at $z=0$, including stellar mass, star formation history, masses of stellar bulge and stellar disc, the radius and height of the stellar disk change by less than $10\%$ between the different realisations, and that magnetic field structures in the disc and the halo are very similar. In contrast, the star formation rate today can vary by a factor of two and the internal morphological structure of the stellar disc can change. The time and orbit of satellite galaxies and their galaxy properties when falling into the main halo are again very similar, but their orbits start to deviate after first pericenter passage. Finally, we show that changing the mass resolution of all matter components in the Auriga model changes galaxy properties significantly more than the intrinsic variability of the model, and that these changes are systematic. This limits detailed comparisons between simulations at different numerical resolutions.
@@
pubyear: 2025pagerange: Quantifying the robustness of the Auriga galaxy formation modelQuantifying the robustness of the Auriga galaxy formation model
@@
Realisation $M_{200\mathrm{c}}$ $M_{*}$ $M_{*,R>50\mathrm{kpc}}$ $\dot{M}_{1\mathrm{Gyr}}$ $V_{\mathrm{c,max}}$ $M_{\mathrm{disc}}$ $M_{\mathrm{bulge}}$ $R_{\mathrm{disc}}$ $H_{\mathrm{disc}}$ $\sigma_{\mathrm{r}}$ $\sigma_{\mathrm{z}}$ $\left|L_{*}\right|$
$[10^{12}\mathrm{M}_{\odot}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[\mathrm{M}_{\odot}/\mathrm{yr}]$ $[\mathrm{km/s}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[\mathrm{kpc}]$ $[\mathrm{kpc}]$ $[\mathrm{km/s}]$ $[\mathrm{km/s}]$ $[\mathrm{10^{18}\mathrm{M}_{\odot}\,kpc\,km/s}]$
@@
[…]
Table 2: Global properties of our galaxy in all $7$ realisations at $z=0$. The columns show from left to right the mass of the halo $M_{200\mathrm{c}}$, that is the mass within a sphere around the potential minimum of the halo where the mean density is 200×200$\times$200 × the critical density of the universe, the stellar mass of the main subhalo $M_{*}$, the stellar mass in the halo at a distance larger than $50\,\mathrm{kpc}$ $M_{*,R>50\mathrm{kpc}}$, the star formation rate in the last Gyr $\dot{M}_{1\mathrm{Gyr}}$, the maximum circular velocity $V_{\mathrm{c,max}}$, the mass of the stellar disc $M_{\mathrm{disc}}$ and bulge $M_{\mathrm{bulge}}$, the scale radius $R_{\mathrm{disc}}$ and height $H_{\mathrm{disc}}$ of the stellar disc, the radial $\sigma_{\mathrm{r}}$ and vertical $\sigma_{\mathrm{z}}$ velocity dispersion in the stellar disc at a radius of $8\,\mathrm{kpc}$, and the total angular momentum of the stellar disc $\left|L_{*}\right|$. The last three rows show the mean value, variance, and the variance relative to the mean for all quantities over all $7$ realisations. Most properties vary by $\lesssim 10\%$. Notably the stellar mass of the halo and the star formation rate at $z=0$ vary by $\approx 30\%$.
@@
1 Introduction
Quantifying the impact of specific physical processes from simulations of a large number of galaxies at reasonable resolution is computationally unfeasible, because every single box is expensive already (Nelson et al., 2019; Villaescusa-Navarro et al., 2021). Therefore, galaxy simulations focusing on the study of physical processes by comparing simulations run with different physics models or parameters usually only simulate a small number of galaxies with a zoom-in approach. This way they focus essentially all resolution elements in one main galaxy of interest. They then simulate the same galaxy multiple times with different physical models, for example to understand the impact of magnetic fields on galaxies (Pakmor et al., 2017; Whittingham et al., 2021, 2023) and the circum-galactic medium (van de Voort et al., 2021), cosmic rays (Buck et al., 2020; Hopkins et al., 2020; Martin-Alvarez et al., 2023; Rodríguez Montero et al., 2024; Bieri & …, 2025), different supermassive black hole models (see, e.g. Irodotou et al., 2022), or the impact of systematic variations of the assembly history of a galaxy (Davies et al., 2021, 2022; Rey et al., 2023; Joshi et al., 2024). Then, however, it becomes critical to understand the robustness of the simulations for individual galaxies.
Despite the fundamental importance for galaxy physics studies, little work has been done focusing on the intrinsic variability of galaxy simulations. Keller et al. (2019) studied the variance in cosmological simulations of dwarf galaxies. They ran two realisations each of the same galaxy with four different models for stellar feedback. They found that the differences between galaxies run with different feedback models were significantly bigger than between pairs of galaxies run with the same feedback model. They then compared each pair of galaxies run with one feedback model to quantify the robustness of each model. They found significant deviations up to a difference in stellar mass of a factor of $2$ in extreme cases for a given feedback model and resolution, and significant difference in the variability for different feedback models.
In this paper we analyse the intrinsic variance of the Auriga galaxy formation model (Grand et al., 2017). Star formation histories and satellite mass functions of our set of $7$ cosmological zoom simulations of the same galaxy have already been shown in Grand et al. (2021) to understand and quantify resolution effects on the satellite population of a Milky Way-like galaxies. Here we present a more general and detailed quantitative analysis of the same $7$ simulations with a focus on the central galaxy at $z=0$.
@@ 2 Simulations
Figure 4: Star formation histories of all seven realisations of Au-6 at resolution L4. They are essentially identical before $z=1$. As the star formation rates decline towards $z=0$, their relative differences increase, up to a deviation of $\sim~{}50\%$ relative to the mean.
@@ 3 Global properties of galaxies and their halo
In Figure 2 we show projections of the dark matter column density convolved with the local dark matter velocity dispersion rendered using a two-dimensional colormap (Springel et al., 2005; Pakmor et al., 2023) at $z=0$. The projections are computed along the $z$-axis of the simulation box, so that the orientation of the main halo and the positions of the satellites are directly comparable. The main halo looks very similar in all realisations with similar orientation and elongation. Moreover, the most massive satellite and many of the smaller satellites are present in all halos, at similar positions. We conclude that the dark matter and satellite halos, including their trajectories, are robustly modelled by our simulations. This holds at least until satellites fall deeply into the main halo, similar to simulations that only evolve dark matter and no baryons (see, e.g. Springel et al., 2008).
To better quantify the difference in the stellar component of the main galaxy we show the star formation histories of all galaxies in the top panel of Figure 4 and the relative deviation from the mean star formation rate in its bottom panel. Here, we include all star particles bound directly to the main halo at $z=0$ (that is, we exclude star particles bound to satellite galaxies). The star formation histories are very similar before $z=1$. After $z=1$ they oscillate around a mean level of $\sim 10\,\mathrm{M_{\odot}/yr}$ until $z\sim 0.4$ and then all decline steeply towards $z=0$. At this time, they have dropped by more than a factor of $5$ to a mean star formation rate of only $2\,\mathrm{M_{\odot}/yr}$, with a significant variance of $1\,\mathrm{M_{\odot}/yr}$. This variance is driven by differences in the gas accretion rate on the galaxy. Even though the relative difference at $z=0$ is quite large, this has almost no effect on the total stellar mass of the galaxy, because it is dominated by stars formed at earlier times when the relative differences between the different simulations are much smaller.
@@ 4 Properties of the galactic discs
Having shown that the stellar discs are similar for the different realisations, we next look at the magnetic fields in the mid-plane of the disc at $z=0$ as a proxy to compare how similar the gas discs are. Magnetic fields affect the gas dynamics and are an important tracer of the state of the gas in both the disc (Pakmor et al., 2014, 2017, 2018; Whittingham et al., 2021, 2023; Pakmor et al., 2024) and the circum-galactic medium (Pakmor et al., 2020; van de Voort et al., 2021). Note that this comparison is only really meaningful because the stellar discs, which dominate the potential in the galaxy and therefore the gas dynamics in the galaxy as well, are so similar.
@@
Figure 10: Star formation histories (top left panel) and stellar mass surface density profiles (top right panel) of the main galaxy for all seven realisations of Au-6 at resolution L4 (purple lines) and L5 (green lines). The shaded regions show interpolated $16\%$ and $84\%$ percentile bands around the mean profile (solid lines). The bottom panels show the same relative to the mean of all simulations at L4. The differences between the two resolution levels in all quantities are significantly larger than the variations between realisations at fixed resolution.
Figure 11: Time evolution of the stellar mass of one matched satellite galaxy of the main galaxy for all seven realisations of Au-6 at the standard resolution L4 (purple lines, the same as in the right panel of Figure 9) and lower resolution L5 (green lines). Similar to the main galaxy the peak stellar mass of the satellite galaxy is lower for the lower resolution simulation. The satellite is destroyed slightly earlier on average in the lower resolution simulation.
5 Satellite galaxies
In addition to the central galaxy, a cosmological zoom-in simulation also contains a number of smaller galaxies, some of them satellites of the main galaxy. The latter are interesting in and of themselves and as a diagnostic for the galaxy evolution model (Simpson et al., 2018; Grand et al., 2021). They also interact with their central host galaxy. They shape its stellar halo and create stellar streams (Monachesi et al., 2016, 2019; Riley et al., 2024; Shipp et al., 2024; Vera-Casanova. et al., 2025), they also influence the disc of the central galaxy, and they contribute to substructures in the disc and inner halo (Gómez et al., 2016, 2017; Simpson et al., 2019; Gargiulo et al., 2019). Therefore, it is crucial to understand how similar the satellites are between realisations when they fall into the halo, to determine how robust their evolution is until their possible disruption.
In addition to looking at satellite galaxies that survive until $z=0$ it is interesting to compare satellites that are destroyed. As an example, we show in Figure 9 the evolution of one satellite that is eventually destroyed. As a representative case, we select the third most massive (by stellar mass) satellite among those disrupted before $z=0$ in R0. We then match it to the other realisations by tracing back the $100$ most bound dark matter particles of each satellite at the time it reaches its peak mass to the initial conditions (see also, Riley et al., 2024). We then compute the centre of mass of these particles, and find the satellite that most closely matches the centre of our reference satellite selected from R0.
@@ 5 Satellite galaxies
Table 3: *
@@
Realisation $M_{200\mathrm{c}}$ $M_{*}$ $M_{*,R>50\mathrm{kpc}}$ $\dot{M}_{1\mathrm{Gyr}}$ $V_{\mathrm{c,max}}$ $M_{\mathrm{disc}}$ $M_{\mathrm{bulge}}$ $R_{\mathrm{disc}}$ $H_{\mathrm{disc}}$ $\sigma_{\mathrm{r}}$ $\sigma_{\mathrm{z}}$ $\left|L_{*}\right|$
$[10^{12}\mathrm{M}_{\odot}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[\mathrm{M}_{\odot}/\mathrm{yr}]$ $[\mathrm{km/s}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[10^{10}\mathrm{M}_{\odot}]$ $[\mathrm{kpc}]$ $[\mathrm{kpc}]$ $[\mathrm{km/s}]$ $[\mathrm{km/s}]$ $[\mathrm{10^{18}\mathrm{M}_{\odot}\,kpc\,km/s}]$
@@
[…]
$\cfrac{\mathrm{mean}-\mathrm{mean_{L4}}}{\mathrm{mean_{L4}}}[\%]$ $-0.8$ $-21.1$ $-39.4$ $-17.5$ $-2.7$ $-14.8$ $-18.8$ $-9.5$ $28.6$ $-2.6$ $-1.1$ $-33.6$

### v2

# Quantifying the intrinsic variability due to randomness of the Auriga galaxy formation model

@@ Abstract
Numerical simulations have become an indispensable tool in astrophysics. To interpret their results, it is critical to understand their intrinsic variability, that is, how much the results change with numerical noise or inherent stochasticity of the physics model. We present a set of seven realisations of high-resolution cosmological zoom-in simulations of a Milky Way-like galaxy with the Auriga galaxy formation model. All realisations share the same initial conditions and code parameters, but draw different random numbers for the inherently stochastic parts of the model. We show that global galaxy properties at $z=0$, including stellar mass, star formation history, masses of stellar bulge and stellar disc, the radius and height of the stellar disk change by less than $10\%$ between the different realisations, and that magnetic field structures in the disc and the halo are very similar. In contrast, the star formation rate today can vary by a factor of two and the internal morphological structure of the stellar disc can change. The time and orbit of satellite galaxies and their galaxy properties when falling into the main halo are again very similar, but their orbits start to deviate after first pericenter passage. Finally, we show that changing the mass resolution of all matter components by a factor of $8$ in the Auriga model changes galaxy properties significantly more than the intrinsic variability of the model, and that these changes are systematic. This limits detailed comparisons between simulations at different numerical resolutions.
@@
pubyear: 2025pagerange: Quantifying the intrinsic variability due to randomness of the Auriga galaxy formation modelQuantifying the intrinsic variability due to randomness of the Auriga galaxy formation model
1 Introduction
Quantifying the impact of specific physical processes from simulations of a large number of galaxies at reasonable resolution is computationally unfeasible, because every single box is expensive already (Nelson et al., 2019; Villaescusa-Navarro et al., 2021). Therefore, galaxy simulations focusing on the study of physical processes by comparing simulations run with different physics models or parameters usually only simulate a small number of galaxies with a zoom-in approach. This way they focus essentially all resolution elements in one main galaxy of interest. They then simulate the same galaxy multiple times with different physical models, for example to understand the impact of magnetic fields on galaxies (Pakmor et al., 2017; Whittingham et al., 2021, 2023) and the circum-galactic medium (van de Voort et al., 2021), cosmic rays (Buck et al., 2020; Hopkins et al., 2020; Martin-Alvarez et al., 2023; Rodríguez Montero et al., 2024; Bieri et al., 2025), different supermassive black hole models (see, e.g. Irodotou et al., 2022), or the impact of systematic variations of the assembly history of a galaxy (Davies et al., 2021, 2022; Rey et al., 2023; Joshi et al., 2024). Then, however, it becomes critical to understand the robustness of the simulations for individual galaxies.
Despite the fundamental importance for galaxy physics studies, little work has been done focusing on the intrinsic variability of galaxy simulations. Keller et al. (2019) studied the intrinsic variance in simulations of isolated dwarf galaxies and cosmological simulations of Milky Way-like galaxies. Using different feedback models they ran pairs of dwarf galaxies with the same model and re-simulated a dwarf galaxy with one model $128$ times. They found a typical relative variation in stellar mass after $1\,\mathrm{Gyr}$ of $5\%$ to $10\%$, with differences up to a factor of $2$ in extreme cases. They also evolved pairs of Milky Way-like galaxies with different feedback models in a fully cosmological setting to $z=0.9$. There they found significantly larger deviations in stellar mass compared to the isolated dwarf galaxies, with persisting differences in stellar mass larger than $10\%$. They also show that the galaxies can appear qualitatively different.
In this paper we analyse the intrinsic variance of the Auriga galaxy formation model (Grand et al., 2017). Star formation histories and satellite mass functions of our set of $7$ cosmological zoom simulations of the same galaxy have already been shown in Grand et al. (2021) to understand and quantify resolution effects on the satellite population of a Milky Way-like galaxies. They found that the stochastic variability at fixed resolution for the Auriga model is significantly smaller than changes introduced by changing the numerical resolution, and that the satellite population is robust to stochastic variations as well. Here we present a more general and detailed quantitative analysis of the same $7$ simulations with a focus on the central galaxy at $z=0$.
@@ 1 Introduction
Realisation $_{200\mathrm{c}} MM_{*}$ $M_{*,R>50\mathrm{kpc}}$ $\dot{M}_{1\mathrm{Gyr}}$ $V_{\mathrm{c,max}}$ $M_{\mathrm{disc}}$ $M_{\mathrm{bulge}}$ $R_{\mathrm{disc}}$ $H_{\mathrm{disc}}$ $\sigma_{\mathrm{r}}$ $\sigma_{\mathrm{z}}$ $\left|L_{*}\right|$
$^{12}\mathrm{M}_{\odot}] [1010$$^{10}\mathrm{M}_{\odot}$] $10^{10}\mathrm{M}_{\odot}] [$$\mathrm{M}_{\odot}/\mathrm{yr}] [$$\mathrm{km/s}] [$$^{10}\mathrm{M}_{\odot}] [1010$$^{10}\mathrm{M}_{\odot}$] $\mathrm{kpc}] [$$\mathrm{kpc}] [$$\mathrm{km/s}] [$$\mathrm{km/s}] [$$\mathrm{10^{18}\mathrm{M}_{\odot}\,kpc\,km/s}]$
@@
$\sigma$ / mean $\%$] $1.0$ $3.2$ $30.6$ $33.6$ $2.2$ $4.0$ $5.5$ $10.7$ $3.1$ $3.9$ $3.1$ $7.2$
Table 1: Global properties of our galaxy in all $7$ realisations at $z=0$. The columns show from left to right the mass of the halo $M_{200\mathrm{c}}$, that is the mass within a sphere around the potential minimum of the halo where the mean density is $200\times$ the critical density of the universe, the stellar mass of the main subhalo $M_{*}$, the stellar mass in the halo at a distance larger than $50\,\mathrm{kpc}$ $M_{*,R>50\mathrm{kpc}}$, the star formation rate in the last Gyr $\dot{M}_{1\mathrm{Gyr}}$, the maximum circular velocity $V_{\mathrm{c,max}}$, the mass of the stellar disc $M_{\mathrm{disc}}$ and bulge $M_{\mathrm{bulge}}$, the scale radius $R_{\mathrm{disc}}$ and height $H_{\mathrm{disc}}$ of the stellar disc, the radial $\sigma_{\mathrm{r}}$ and vertical $\sigma_{\mathrm{z}}$ velocity dispersion in the stellar disc at a radius of $8\,\mathrm{kpc}$, and the total angular momentum of the stellar disc $\left|L_{*}\right|$. The last three rows show the mean value, variance, and the variance relative to the mean for all quantities over all $7$ realisations. Most properties vary by $\lesssim 10\%$. Notably the stellar mass of the halo and the star formation rate at $z=0$ vary by $\approx 30\%$.
@@ 2 Simulations
Figure 4: Star formation histories of all seven realisations of Au-6 at standard resolution L4 (top panel). They are essentially identical before $z=1$. As the star formation rates decline towards $z=0$, their relative differences (middle panel) increase, up to a deviation of $\sim~50\%$ relative to the mean. The total stellar mass deviates only on the level of a few percent over the whole evolution (bottom panel).
@@
In addition to randomness introduced by the physics model, we randomize the centre of the domain in every domain decomposition to decorrelate and therefore reduce force errors (Springel et al., 2021). The refinement also introduces randomness when deciding where to exactly introduce a new mesh-generating point when a cell is split (Springel, 2010).
@@
Figure 6: Metallicity profiles of stars (left panels) and gas (right panels) in the disc at $z=0$ in cylindrical coordinates for all seven realisations of Au-6 at the standard resolution L4. In the inner $15\,\mathrm{kpc}$ the stellar metallicity profile is essentially identical, at larger radii it varies from $0.5\,\mathrm{Z_{\odot}}$ to $\mathrm{Z_{\odot}}$. The gas metallicity shows slightly larger scatter in the inner parts and slightly less scatter in the outer parts, compared to the stellar metallicity profiles.
@@ 3 Global properties of galaxies and their halo
In Figure 2 we show projections of the dark matter column density convolved with the local dark matter velocity dispersion rendered using a two-dimensional colormap (Springel et al., 2005; Pakmor et al., 2023) at $z=0$. The projections are computed along the $z$-axis of the simulation box, so that the orientation of the main halo and the positions of the satellites are directly comparable. The main halo looks very similar in all realisations with similar orientation and elongation. Moreover, the most massive satellite and many of the smaller satellites are present in all halos, at similar positions. We conclude that the dark matter and satellite halos, including their trajectories, are robustly modelled by our simulations. This holds at least until satellites fall deeply into the main halo, similar to simulations that only evolve dark matter and no baryons (see, e.g. Springel et al., 2008). The deviations in the positions of smaller halos relative to the main halo are caused by errors in the force calculation. Those are directly exposed here, because we add a random shift vector to the global domain at every domain decomposition to decorrelate force errors over time. This reduces the overall force error (Springel et al., 2021), and also leads to different force errors in the different realisations.
To better quantify the difference in the stellar component of the main galaxy we show the star formation histories of all galaxies in the top panel of Figure 4, the relative deviation from the mean star formation rate in its middle panel, and the time evolution of the relative deviation of the stellar mass from the mean stellar mass in the bottom panel. Here, we include all star particles bound directly to the main halo at $z=0$ (that is, we exclude star particles bound to satellite galaxies). The star formation histories are essentially identical before $z=2$ and very similar before $z=1$. The initial deviation might be connected to either a significant merger just before $z=2$ or the formation of a disc. After $z=1$ they oscillate around a mean level of $\sim 10\,\mathrm{M_{\odot}/yr}$ until $z\sim 0.4$ and then all decline steeply towards $z=0$. At this time, they have dropped by more than a factor of $5$ to a mean star formation rate of only $2\,\mathrm{M_{\odot}/yr}$, with a significant variance of $1\,\mathrm{M_{\odot}/yr}$. This variance is driven by differences in the gas accretion rate on the galaxy. Even though the relative difference at $z=0$ is quite large, this has almost no effect on the total stellar mass of the galaxy, because it is dominated by stars formed at earlier times when the relative differences between the different simulations are much smaller. The total stellar mass only deviates on the level of a few per cent at all times.
@@ 4 Properties of the galactic discs
We show stellar metallicity profiles and their relative deviation from the mean profile in the left panels of Figure 6. In the inner $15\,\mathrm{kpc}$ the stellar metallicity profile only deviates by a few per cent. At larger radii the deviation is significantly larger. In the very outskirts of the stellar disc at $30\,\mathrm{kpc}$ the stellar metallicity varies from $0.5\,\mathrm{Z_{\odot}}$ to $\mathrm{Z_{\odot}}$. The deviation from the mean stellar metallicity profile (see lower left panel of Figure 6) in the outskirts strongly correlates with the deviation from the mean stellar surface density profile (see lower left panel of Figure 5). Higher stellar surface density leads to higher stellar metallicity and vice versa.
Having shown that the stellar discs are similar for the different realisations, we next look at gas metallicity profiles at $z=0$ in the right panels of Figure 6. The gas metallicity profiles have slightly larger scatter in the inner parts of the disc compared to the stellar metallicity profiles, and slightly smaller scatter in the outer parts. The scatter is still below $20\%$ at all radii.
We then look at the magnetic fields in the mid-plane of the disc at $z=0$ as a proxy to compare how similar the gas discs are. Magnetic fields affect the gas dynamics and are an important tracer of the state of the gas in both the disc (Pakmor et al., 2014, 2017, 2018; Whittingham et al., 2021, 2023; Pakmor et al., 2024) and the circum-galactic medium (Pakmor et al., 2020; van de Voort et al., 2021). Note that this comparison is only really meaningful because the stellar discs, which dominate the potential in the galaxy and therefore the gas dynamics in the galaxy as well, are so similar.
@@
5 Satellite galaxies
In addition to the central galaxy, a cosmological zoom-in simulation also contains a number of smaller galaxies, some of them satellites of the main galaxy. The latter are interesting in and of themselves and as a diagnostic for the galaxy evolution model (Simpson et al., 2018; Grand et al., 2021). They also interact with their central host galaxy. They shape its stellar halo and create stellar streams (Monachesi et al., 2016, 2019; Riley et al., 2025; Shipp et al., 2025; Vera-Casanova. et al., 2025), they also influence the disc of the central galaxy, and they contribute to substructures in the disc and inner halo (Gómez et al., 2016, 2017; Simpson et al., 2019; Gargiulo et al., 2019). Therefore, it is crucial to understand how similar the satellites are between realisations when they fall into the halo, to determine how robust their evolution is until their possible disruption.
In addition to looking at satellite galaxies that survive until $z=0$ it is interesting to compare satellites that are destroyed. As an example, we show in Figure 10 the evolution of one satellite that is eventually destroyed. As a representative case, we select the third most massive (by stellar mass) satellite among those disrupted before $z=0$ in R0. We then match it to the other realisations by tracing back the $100$ most bound dark matter particles of each satellite at the time it reaches its peak mass to the initial conditions (see also, Riley et al., 2025). We then compute the centre of mass of these particles, and find the satellite that most closely matches the centre of our reference satellite selected from R0.
@@ 5 Satellite galaxies
Figure 11: Star formation histories (top left panel) and stellar mass surface density profiles (top right panel) of the main galaxy for all seven realisations of Au-6 at standard resolution L4 (purple lines) and 88 times lower resolution L5 (green lines). The shaded regions show interpolated $16\%$ and $84\%$ percentile bands around the mean profile (solid lines). The bottom panels show the same relative to the mean of all simulations at L4. The differences between the two resolution levels in all quantities are significantly larger than the variations between realisations at fixed resolution.
Figure 12: Time evolution of the stellar mass of one matched satellite galaxy of the main galaxy for all seven realisations of Au-6 at the standard resolution L4 (purple lines, the same as in the right panel of Figure 10) and 88 times lower resolution L5 (green lines). Similar to the main galaxy the peak stellar mass of the satellite galaxy is lower for the lower resolution simulation. The satellite is destroyed slightly earlier on average in the lower resolution simulation.
@@
Realisation $_{200\mathrm{c}} MM_{*}$ $M_{*,R>50\mathrm{kpc}}$ $\dot{M}_{1\mathrm{Gyr}}$ $V_{\mathrm{c,max}}$ $M_{\mathrm{disc}}$ $M_{\mathrm{bulge}}$ $R_{\mathrm{disc}}$ $H_{\mathrm{disc}}$ $\sigma_{\mathrm{r}}$ $\sigma_{\mathrm{z}}$ $\left|L_{*}\right|$
$^{12}\mathrm{M}_{\odot}] [1010$$^{10}\mathrm{M}_{\odot}$] $10^{10}\mathrm{M}_{\odot}] [$$\mathrm{M}_{\odot}/\mathrm{yr}] [$$\mathrm{km/s}] [$$^{10}\mathrm{M}_{\odot}] [1010$$^{10}\mathrm{M}_{\odot}$] $\mathrm{kpc}] [$$\mathrm{kpc}] [$$\mathrm{km/s}] [$$\mathrm{km/s}] [$$\mathrm{10^{18}\mathrm{M}_{\odot}\,kpc\,km/s}]$
@@
$\sigma$ / mean $\%$] $0.9$ $7.3$ $40.6$ $58.3$ $1.8$ $2.7$ $6.7$ $11.4$ $12.8$ $4.7$ $3.9$ $6.6$
$\cfrac{\mathrm{mean}-\mathrm{mean_{L4}}}{\mathrm{mean_{L4}}}[\%]$ 0.8-0.8 21.1-21.1 39.4-39.4 17.5-17.5 2.7-2.7 14.8-14.8 18.8-18.8 9.5-9.5 $28.6$ 2.6-2.6 1.1-1.1 33.6-33.6